# Electroweak Monopole-Antimonopole Pair Production at LHC

### v4

@@ IV Electroweak Monopole Production from p-p Fireball at LHC
This means that the energy condition to produce the monopole mass no more than 7 TeV forces the monopole production at LHC to stop at 119.4 GeV, much higher that the Ginzburg temperature. The effective potential at $T_{min}^{LHC}$ is also shown in Fig. 3 for comparison.
If so, at LHC the electroweak monopole production starts at $T_{c}$ around 146.7 GeV and stops at $T_{min}^{LHC}$ around 119.4 GeV, not at the Ginzburg temperature. In average the monopole production temperature at LHC is given by
@@
It is generally believed that in the Kibble-Zurek mechanism we are supposed to have one monopole per one correlation volume. This assumotion, however, may have a critical defect. This is because the correlation length is fixed by the electroweak scale but the monopole mass is given by the $1/\alpha$ times bigger than the electroweak scale, so that the energy in one correlation volume may not be enough to make up the monopole mass. A natural way to cure this defect is to enlarge the correlation length $\xi=1/\bar{M}_{H}$ to $\bar{\xi}$,
@@
Unfortunately, the above argument has a critical defect, because here we have implicitly assumed that the temperature of the p-p fireball would be high enough to produce the monopole thermally (assuming that the 14 TeV p-p fireball energy would produce the high temperature necessary for LHC to produce the monopole thermally). But this is not obvious, because the temperature and energy of the fireball are two different things. As we have argued, for LHC to produce the monopole thermally, the temperature of the fireball must be no less than 119.4 GeV. And if the temperature of the fireball higher than this, the above conclusion becomes valid. But is this really so? How can we guarantee this? This is a non-trivial question.
A straightforward way to answer this question is to measure the temperature of the p-p firebal at LHC. And actually the ALICE group has measured the temperature of the p-p fireball of the center of mass energy 2.76 TeV at LHC, and found that the temperature is around 297 MeV, which is roughly $10^{-4}$ times smaller than the center of mass energy of the p-p fireball [27]. The ALICE result was unexpected, but clearly shows that the temperature of LHC p-p fireball is much less than the temperature to produce the monopole thermally, even with the future FCC energy upgrade.
This tells that here again, the thermal production of the electroweak monopole at LHC is practically impossible, although it is logically possible. This teaches us an important lesson. If the monopole production mechanism at LHC is the thermal production, it is vertually impossible for LHC to produce the monopole, simply because the temperature of the p-p fireball is too low. In this case the only way to detect the electroweak monopole is to look for the remnant monopoles produced in the earluy universe.
This also tells that the only probable way for LHC to prooduce the monopole is by the Drell-Yan process and/or the Schwinger mechanism. Obviously, the present 14 TeV LHC could produce the monopole if the mass becomes less than 7 TeV, if the monopole production mechanism becomes the Drell-Yan process and/or the Schwinger mechanism. But as we have pointed out, in this case there is a possibility for LHC to produce the monopole in the form of the monopolium bound states, even when the monopole mass becomes 11 TeV. This is because the monopolium binding energy could reduse the monopolium mass below 14 TeV even when the mass of the monopole-antimonopole pair becomes 22 TeV. This necessiate us to study the monopolium bound state, which we discuss in the following.
@@ V Electroweak Monopolium: Atomic Model
The monopolium itself has been studied before for various reasons [28, 29, 30]. For example, Nambu proposed to view the monopolium as a model of strong interaction, because the magnetic coupling of the monopole-antionopole pair drastically increases the energy spectrum to the order of hundred MeV [28]. But here we are interested in the electroweak monopolium made of the Cho-Maison monopole pair.
@@
Notice, however, that these results are totally unreasonable. First, the above Rydburg energy is too big (roughly 2,346 times bigger than the monopole-antimonopole mass), which means that the monopole mass simply can not provide the binding energy necessary to make the monopolium. Second, the monopolium size (i.e., the Bohr radius) is $10^{-8}$ times smaller than the monopole size.
This could mean the following. First, the monopolium quickly annihilates itself, long before it reaches to the ground state. In fact (50) suggests that the monopolium could exist only at highly excited levels, probably with $n\geq 49$, because below this level the monopole-antimonopole mass can not provide the necessary binding energy. Or else, this could mean that the naive atomic description of the monopolium is not acceptable. In this case we have to find a more realistic model of the monopolium.
The reason for the above results originates from the fact that the magnetic coupling of the monopole is too strong, which makes the monipolium size too small. We could make the atomic model of the monopolium more realistic by improving on this point. For example, we could modify the magnetic Coulombic potential at short distance for $r\leq 1/M_{W}$, to accommodate the fact that the monopole has a finite size $1/M_{W}$.
@@
VI A Realistic Electroweak Monopolium
To have a realistic model of the electroweak monopolium, we have to remember that intuitively the size of the electroweak monopolium can not be smaller than $2/M_{W}$, considering the fact that the monopole (and antiminopole) has the size of $1/M_{W}$. This means that at short distance the singularity of the magnetic potential in (49) should be regularized.
@@ VI A Realistic Electroweak Monopolium
$V=\left\{\begin{array}[]{rcl}\dfrac{2\pi}{e^{2}r_{0}}\Big{(}\dfrac{r^{2}}{r_{0}^{2}}-3\Big{)},&~{}~{}~{}r<r_{0},\\ {}\hfil\\ -\dfrac{4\pi}{e^{2}}\dfrac{1}{r},{}{}{}{}{}{}{}{}&~{}~{}~{}r\geq r_{0},\end{array}\right.$ (55)
@@
The first step is the usual separation of variables for wave functions in the regions $r<r_{0}$ (region I) and $r>r_{0}$ (region II) as
$\Psi_{\rm{I,II}}(r,\varphi,\theta)\equiv\frac{R_{\rm{I,II}}(r)}{r}Y_{l}^{m}(\varphi,\theta)\,,$ (56)
where $Y_{l}^{m}(\varphi,\theta)$ are the spherical harmonics. Next, we adopt dimensionless coordinates
@@
In the region I ($x<x_{0}$), we need to satisfy the regularity condition at the origin $x=0$, which force us to exclude $U(a,b,z)$ as it has a singularity at $z=0$ in contrast to $M(a,b,z)$, which is an entire function in $z$. Hence, the general solution in the first region that is regular at the origin is given by
@@
×$\times_{1}F_{1}\Big{(}\begin{matrix}\frac{l}{2}+\frac{3}{4}+\frac{x_{0}^{3/2}\epsilon-3\sqrt{x_{0}}}{4}\\ l+\frac{3}{2}\end{matrix}\Bigr{|}\tfrac{x^{2}}{x_{0}^{3/2}}\Big{)},$ (62)
where $N_{\rm I}$ is an arbitrary constant at the moment.
In the region II ($x>x_{0}$), we need to impose boundary condition at $x\to\infty$ so that the resulting wave equation is square-integrable. This force us to discard $M(a,b,z)$ that behaves as $M(a,b,z)\sim\Gamma(b)\exp(z)~{}z^{a-b}/\Gamma(a)$ for large $z$ (in a certain wedge of the complex plane). Notice that this asymptotic behaviour holds only if $a$ is not a negative integer (and hence $\Gamma(a)$ blows up), which is exactly the case for the hydrogen atom. For our modified potential, however, the spectrum differs and hence $a\not=-k$, where $k$ is a positive integer.
@@
as $x$ goes to infinity. Hence, the general solution in the second region has the following form,
@@
×$\times U\Big{(}\begin{matrix}l+1-\frac{1}{\sqrt{\epsilon}}\\ 2l+2\end{matrix}\Bigr{|}2\sqrt{\epsilon}x\Big{)},$ (64)
where $N_{\rm II}$ is again an arbitrary constant.
Figure 8: The energy spectrum $\sqrt{\epsilon}$ of the monopolium with $l=0$ fixed by (65), as function of $\sqrt{x_{0}}$. The dashed lines represent harmonic oscillator energy levels given by (67). The horizontal dotted lines are the energy levels of the hydrogen atom given by (66).
The remaining task is to formulate sewing condition that ensures $C_{1}$ continuity of the wavefunction at the point $x=x_{0}$. The condition that $R_{\rm I}(x_{0})=R_{\rm II}(x_{0})$ fixes, say, $N_{\rm II}$ in terms of $N_{\rm I}$, which can be in turn fixed by square integrability condition. The other condition $R_{\rm I}^{\prime}(x_{0})=R_{\rm II}^{\prime}(x_{0})$ then boils down to a transcendental equation for the binding energy,
@@
With the two conditions we have the smooth radial wave function of the monopolium. This is shown in Fig. 7.
Notice that the above equation determines the energy spectrum of the monopolium, so that it is very important to solve (8). Unfortunately, this condition can only be solved numerically, and we display the energy spectrum in Fig. 8 (for $l=0$).
@@
The value of $x_{0}$ we are interested is given by $r_{0}\sim 1/M_{W}$, or
@@
For such a high values of $x_{0}$, the binding energies are almost indistinguishable from binding energies of harmonic oscillator.
@@
where we have put $M\simeq 11.0~{}\text{TeV}$. Compared with (50), this is roughly $3\times 10^{-4}$ times reduction of the naive estimate of the binding energy given by hydrogen-like atomic model of the monopolium. In particular, this tells that the monopole-antimonopole pair with mass 22.0 TeV makes the monopolium bound state of mass of 5.2 TeV at the ground state. This is precisely what we have hoped for.
@@
Notice that in the above analysis we have assumed the monopole mass to be 11.0 TeV to have the monopolium mass 5.2 TeV. And (69) tells that the monopolium mass becomes 3.3 TeV when the monopole mass becomes 7 TeV. Perhaps more importantly, this tells that the maximum mass of the monopole that the present 14 TeV LHC can produce the monopolium of mass 14 TeV becomes 29.8 TeV. This strongly implies that the present LHC could produce the electroweak monopole in the form of the monopolium bound states for any reasonable monopole mass (as far as the mass remains less than 29.8 TeV), if the monopole production mechanisn becomes the Drell-Yan process.
@@
has been discussed before [30]. Despite the obvious difference between our potential (55) and this, we notice that the characteristic features of the two monopoliums are quite similar. In particular, in both cases the monopolium masses are slightly below the half of the monopole’s mass and the radiuses are of the order of $10^{-18}cm$. This implies that our results discussed above are reliable.
@@ VII Discussion
An urgent task at LHC is to discover the electroweak monopole. Because of the unique charactistics of the electroweak monopole, the MoEDAL detector may have no difficulty to detect the monopole, if LHC could produce it. But the problem is that intuitively the present 14 TeV LHC may not be able to produce the monopole, if the monopole mass becomes bigger than 7 TeV. And indeed, if we suppose the monopole mass is $1/\alpha$ times the W boson mass, LHC cannot produce it. In this case the detection of the monopole at LHC may become hopeless, and we may have to look for the remnant electroweak monopoles which could have been produced in the early universe after the electroweak phase transition [15, 18].
In this paper we have shown that the question if the present LHC could produce the monopole or not critically depends on the monopole production mechanism at LHC. Our work in this paper tells the followings. First, if the monopole production mechanism at LHC is the topological thermal fluctuation of the Higgs vacuum, there is practically no hope that it could produce the monopole, simply because the temperature of the p-p fireball is too low to produce the monopole thermally. This is unexpected, but if the recent ALICE measurement of the temperature of the p-p fireball is trustable, this conclusion is unavoidable.
Second, if the monopole production mechanism at LHC is the Drell-Yan process, the present 14 TeV LHC could produce the monopole if the monopole mass becomes less than 7 TeV. This was expected. What is new is that, more probably LHC could produce the monopole in the form of the monopolium bound states for any reasonable mass of the monopole, as far as the mass does not exceed 29,7 TeV. This is because the monopolium binding energy greatly reduces the mass of the monopolium.
This tells that the present LHC could circumbent the energy constraint and produce the electroweak monopole even when the monopole mass becomes bigger than 7 TeV. This is remarkable.
@@
To summarize, our results tells that a most probable way for LHC to produce the monopole is the Drell-Yan process, and the 14 TeV LHC could more likely produce the monopole in the form of the monopolium bound states than as individual monopoles. In this case the most realistic way to detect the monopole at LHC is to detect the two photon decay modes of the monopoliums. And in principle ATLAS and CMS could do that.
But there is the possibility that the LHC (with the next energy upgrading) may not produce the monopole signal at all. This could mean that the monopole production mechanism at LHC is probably not the Drell-Yan process. In this case the only way to detect the electroweak monopole could be to look for the remnant monopoles produced during the early universe, with the “cosmic” MoEDAL. This makes the detection of the remnant monopoles produced during the early universe very important.

### v5

@@ IV Topological Monopole Production by thermal fluctuation of Higgs vacuum at LHC
This means that the energy condition to produce the monopole mass no more than 7 TeV forces the monopole production at LHC to stop at 119.4 GeV, much higher that the Ginzburg temperature. The effective potential at $T_{min}^{LHC}$ is also shown in Fig. 3 in green line for comparison.
If so, at LHC the electroweak monopole production starts at $T_{c}$ around 146.7 GeV and stops at $T_{min}^{LHC}$ around 119.4 GeV, not at the Ginzburg temperature around 57.5 GeV. In average the monopole production temperature at LHC is given by
@@
It is generally believed that in the Kibble-Zurek mechanism we are supposed to have one monopole per one correlation volume. This assumotion, however, may have a critical defect. This is because the correlation length is fixed by the electroweak scale but the monopole mass is given by $1/\alpha$ times bigger than the electroweak scale, so that the energy in one correlation volume may not be enough to make up the monopole mass. A natural way to cure this defect is to enlarge the correlation length $\xi=1/\bar{M}_{H}$ to $\bar{\xi}$,
@@
Unfortunately, the above argument has a critical defect, because here we have implicitly assumed that the temperature of the p-p fireball would be high enough to produce the monopole thermally (assuming that the 14 TeV p-p fireball energy would produce the high temperature necessary for LHC to produce the monopole thermally). However, this is not obvious, because the temperature and energy of the fireball are two different things. As we have argued, for LHC to produce the monopole thermally, the temperature of the fireball must be no less than 119.4 GeV. And if the temperature of the fireball higher than this, the above conclusion becomes valid. But is this really so? This is a non-trivial question.
A straightforward way to answer this question is to measure the temperature of the p-p firebal at LHC. Fortunately the ALICE group has actually measured the temperature of the p-p fireball of the center of mass energy 2.76 TeV at LHC, and found that the temperature is around 297 MeV, which is roughly $10^{-4}$ times smaller than the center of mass energy of the p-p fireball [30]. The ALICE result was unexpected, but clearly shows that the temperature of LHC p-p fireball is much less than the temperature needed to produce the monopole thermally, even with the future FCC energy.
This tells that here again, the thermal production of the electroweak monopole at LHC is practically impossible, although it is logically possible. This teaches us an important lesson. If the monopole production mechanism at LHC is the thermal production, it is virtually impossible for LHC to produce the monopole, simply because the temperature of the p-p fireball is too low. In this case the only way to detect the electroweak monopole is to look for the remnant monopoles produced in the earluy universe.
This also seems to suggest that the Drell-Yan process and/or the Schwinger mechanism might be the only probable way for LHC to produce the monopole. Obviously, the present 14 TeV LHC could produce the monopole when the mass becomes less than 7 TeV, if the monopole production mechanism becomes the Drell-Yan process and/or the Schwinger mechanism.
On the other hand, there may be another logical possibility for LHC to produce the monopole, the monopole production in the form of the monopolium bound states, as we have pointed out. An important feature of this mechanism is that the binding energy of the monopolium could reduce the mass below 14 TeV even when the mass of the monopole-antimonopole pair becomes 22 TeV. In this case LHC could produce the monopolium even when the monopole mass becomes 11 TeV or heavier. This necessiate us to study the monopolium bound state in more detail, which we discuss in the following.
@@ V Electroweak Monopolium: Atomic Model
The monopolium itself has been studied before for various reasons [31, 32, 33]. For example, Nambu proposed to view the monopolium as a model of strong interaction, because the magnetic coupling of the monopole-antionopole pair drastically increases the energy spectrum of the monopole pair to the order of hundred MeV [31]. But here we are interested in the electroweak monopolium made of the Cho-Maison monopole pair.
@@
Notice, however, that these results are totally unreasonable. First, the above Rydburg energy is too big (roughly 2,346 times bigger than the monopole-antimonopole mass), which means that the monopole mass simply can not provide the binding energy necessary to make the monopolium. Second, the monopolium size (i.e., the Bohr radius) is $10^{-8}$ times smaller than the monopole size, which is absurd.
Figure 6: The monopole production mechanismin the form of the monopolium at LHC.
This could mean the following. First, the monopolium quickly annihilates itself, long before it decays to the ground state. In fact (50) suggests that the monopolium could exist only at highly excited levels, probably with $n\geq 49$, because below this level the monopole-antimonopole mass can not provide the necessary binding energy. Or else, this could mean that the naive atomic description of the monopolium is not acceptable. In this case we have to find a more realistic model of the monopolium.
The reason for the above results originates from the fact that the magnetic coupling of the monopole is too strong, which makes the monopolium size too small. We could make the atomic model of the monopolium more realistic by improving on this point. For example, we could modify the magnetic Coulombic potential at short distance for $r\leq 1/M_{W}$, to accommodate the fact that the monopole has a finite size $1/M_{W}$.
@@
VI Electroweak Monopole Production at LHC via Monopolium
Now, we are ready to discuss the new monopole production mechanism in the form of the monopolium, which could allow us to prove the existence of the electroweak monopole at LHC even when the monopole mass exceeds 11 TeV. This is schimatically shown in Fig. 6. Notice that LHC could produce the monopole with this mechanism, when the monopole production mechanism at LHC becomes the Drell-Yan process and/or the Schwinger mechanism but the energy condition prevent the production of the monopole when the monopole mass becomes bigger than 7 TeV.
In fact the first stage of this mechanism is exactly like the Drell-Yan process which produces the monopole-antimonopole pair. But when the monopole mass exceeds 7 TeV, the pair can not be materialized and could exist only vertually. So, the Drell-Yan process becomes “virtual”, and the final state of this mechanism becomes not the monopole-antimonopole pair but the decay modes of the monopolium.
We could have a similar situation with the Schwinger mechanism. Here again the energy condition of LHC could force the monopole-antimonopole pair virtual, and the final state could become the decay modes of the monopolium. In this sence this monopole production mechanism could be thought as the virtual Drell-Yan process or the virtual Schwinger mechanism. On the other hand, we emphasize that the final state of this monopole production mechnism is totally different from the Drell-Yan process and/or the Schwinger mechanism.
To show that this mechanism could indeed prove the existence of the electroweak monopole at LHC even when the monopole mass exceeds 11 TeV, we must have a realistic model of the monopolium whose binding enerfy could reduce the monopolum mass below 14 TeV. For this we have to remember that intuitively the size of the electroweak monopolium can not be smaller than $2/M_{W}$, considering the fact that the monopole has the size of $1/M_{W}$. This means that at short distance the singularity of the magnetic potential in (49) should be regularized.
@@ VI Electroweak Monopole Production at LHC via Monopolium
$V=\left\{\begin{array}[]{rcl}\dfrac{2\pi}{e^{2}r_{0}}\Big{(}\dfrac{r^{2}}{r_{0}^{2}}-3\Big{)},&~~~r<r_{0},\\ \\ -\dfrac{4\pi}{e^{2}}\dfrac{1}{r},&~~~r\geq r_{0},\end{array}\right$. (55)
@@
The first step to do this is the usual separation of variables of the wave function in the regions $r<r_{0}$ (the region I) and $r>r_{0}$ (the region II) as
$\Psi_{\rm{I,II}}(r,\theta,\varphi)\equiv\frac{R_{\rm{I,II}}(r)}{r}Y_{l}^{m}(\theta,\varphi),$ (56)
where $Y_{l}^{m}(\theta,\varphi)$ are the spherical harmonics. Next, we adopt dimensionless coordinates
@@
In the region I ($x<x_{0}$), we need to satisfy the regularity condition at the origin $x=0$, which forces us to exclude $U(a,b,z)$ as it has a singularity at $z=0$ in contrast to $M(a,b,z)$, which is an entire function in $z$. Hence, the general solution in the first region that is regular at the origin is given by
@@
$\times_{1}F_{1}\Big{(}\begin{matrix}\frac{l}{2}+\frac{3}{4}+\frac{x_{0}^{3/2}\epsilon-3\sqrt{x_{0}}}{4}\\ l+\frac{3}{2}\end{matrix}\Bigr{|}\tfrac{x^{2}}{x_{0}^{3/2}}\Big{)},$ (62)
where $N_{\rm I}$ is an arbitrary constant at the moment which should be determined later.
In the region II ($x>x_{0}$), we need to impose boundary condition at $x\to\infty$ so that the resulting wave equation is square-integrable. This forces us to discard $M(a,b,z)$ that behaves as $M(a,b,z)\sim\Gamma(b)\exp(z)~z^{a-b}/\Gamma(a)$ for large $z$ (in a certain wedge of the complex plane). Notice that this asymptotic behaviour holds only if $a$ is not a negative integer (and hence $\Gamma(a)$ blows up), which is exactly the case for the hydrogen atom. For our modified potential, however, the spectrum differs and hence $a\neq-k$, where $k$ is a positive integer.
@@
as $x$ goes to infinity. Hence, the general solution in the region II has the following form,
@@
$\times U\Big{(}\begin{matrix}l+1-\frac{1}{\sqrt{\epsilon}}\\ 2l+2\end{matrix}\Bigr{|}2\sqrt{\epsilon}x\Big{)},$ (64)
where $N_{\rm II}$ is again an arbitrary constant for the moment.
The remaining task is to formulate sewing condition that ensures $C_{1}$ continuity of the wavefunction at the point $x=x_{0}$. The condition that $R_{\rm I}(x_{0})=R_{\rm II}(x_{0})$ fixes $N_{\rm II}$ in terms of $N_{\rm I}$, which can be in turn fixed by square integrability condition. The other condition $R_{\rm I}^{\prime}(x_{0})=R_{\rm II}^{\prime}(x_{0})$ then boils down to the following transcendental equation for the binding energy,
@@
With the two conditions we can obtain the smooth radial wave function and the energy spectrum of the monopolium.
The radial wave function for n=1,2,,5n=1,2,...,5 is shown in Fig. 8. However, the condition (65) which determines the energy spectrum of the monopolium is not easy to solve, and it can only be solved numerically. Fortunately we could solve it, and the resulting energy spectrum is displayed in Fig. 9 (for $l=0$).
@@
The value of $x_{0}$ we are interested in is given by $r_{0}\sim 1/M_{W}$, or
@@
For such a high values of $x_{0}$, the binding energies are almost indistinguishable from the binding energies of harmonic oscillator.
@@
where we have put $M\simeq 11.0~\text{TeV}$. Compared with (50), this is roughly $3\times 10^{-4}$ times reduction of the naive estimate of the binding energy given by hydrogen-like atomic model of the monopolium. In particular, this tells that the monopole-antimonopole pair with mass 22 TeV makes the monopolium bound state of mass of 5.2 TeV at the ground state. This is precisely what we have hoped for.
Figure 9: The energy spectrum $\sqrt{\epsilon}$ of the monopolium with $l=0$ fixed by (65), as function of $\sqrt{x_{0}}$. The horizontal dotted lines are the energy levels of the hydrogen atom given by (66), and the dashed lines represent harmonic oscillator energy levels given by (67).
@@
Notice that in the above analysis we have assumed the monopole mass to be 11 TeV to have the monopolium mass 5.2 TeV. And (69) tells that the monopolium mass becomes 3.3 TeV when the monopole mass becomes 7 TeV. Perhaps more importantly, this tells that the maximum mass of the monopole that the present 14 TeV LHC can produce the monopolium of mass 14 TeV becomes 29.8 TeV. This strongly implies that the present LHC could produce the electroweak monopole in the form of the monopolium bound states for any reasonable monopole mass (as far as the mass remains less than 29.8 TeV), if the monopole production mechanisn becomes the Drell-Yan process.
@@
has been discussed before [33]. Despite the obvious difference between our potential (55) and this, we notice that the characteristic features of the two monopoliums are quite similar. In particular, in both cases the monopolium masses are slightly below the half of the monopole mass and the radius are of the order of $10^{-18}cm$. This implies that our results discussed above are reliable.
@@ VII Discussion
An urgent task at LHC is to discover the electroweak monopole which exists within the standard model. Because of the unique charactistics of the electroweak monopole, the MoEDAL detector may have no difficulty to detect the monopole, if LHC could produce it. But the problem is that intuitively the present 14 TeV LHC may not be able to produce the monopole, if the monopole mass becomes bigger than 7 TeV. And indeed, if we suppose the monopole mass is $1/\alpha$ times the W boson mass, the present LHC cannot produce it. In this case the detection of the monopole at LHC may become hopeless, and we may have to look for the remnant electroweak monopoles which could have been produced in the early universe after the electroweak phase transition [8, 20].
In this paper we have shown that the question if the present LHC could produce the monopole or not critically depends on the monopole production mechanism at LHC. Our work in this paper strongly implies the followings. First, if the monopole production mechanism at LHC is the topological thermal fluctuation of the Higgs vacuum, there is practically no hope that it could produce the monopole, simply because the temperature of the p-p fireball is too low to produce the monopole thermally. And this would be the case even with the new FCC. This is unexpected, but if the recent ALICE measurement of the temperature of the p-p fireball is trustable, this conclusion is unavoidable.
Second, if the monopole production mechanism at LHC is the Drell-Yan process, the present 14 TeV LHC could produce the monopole only if the monopole mass becomes less than 7 TeV. This was expected.
In this paper we discussed a new monopole production mechanism, the monopole production in the form of the monopolium bound state, which could allow LHC to produce the electroweak monopole for any reasonable mass of the monopole, as far as the mass does not exceed 29,7 TeV. This is because the monopolium binding energy greatly reduces the mass of the monopolium. This tells that the present LHC could circumbent the energy constraint and produce the electroweak monopole in the form of the monopolium bound state, even when the monopole mass becomes bigger than 7 TeV. In this case we MoEDAL, ATLAS, and CMS could detect the decay modes of the monopolium and thus confirm the existence of the electroweak monopole at LHC. This is remarkable.
@@
To summarize, our results tells that a most probable way for LHC to produce the monopole is the monopole production in the form of the monopolium, and the 14 TeV LHC could more likely produce the monopole in the form of the monopolium bound states rather than as individual monopoles. In this case the most realistic way to detect the monopole at LHC is to detect the two photon decay modes of the monopoliums. And in principle MoEDAL, ATLAS, and CMS could do that.
Finally, it should also be emphasized that there is the possibility that the LHC may not produce the monopole signal at all. This is because neither the Drell-Yan process nor the Schwinger mechanism could turn out to be the monopole production mechanism at LHC. In this case the only way to detect the electroweak monopole could be to look for the remnant monopoles produced during the early universe, with the “cosmic” MoEDAL. This makes the detection of the remnant monopoles produced during the early universe very important.